# Sotopia-RL: Reward Design for Social Intelligence

### v1

@@ Abstract
Social intelligence has become a critical capability for large language models (LLMs), enabling them to engage effectively in real-world social tasks such as accommodation, persuasion, collaboration, and negotiation. Reinforcement learning (RL) is a natural fit for training socially intelligent agents because it allows models to learn sophisticated strategies directly through social interactions. However, social interactions have two key characteristics that set barriers for RL training: (1) partial observability, where utterances have indirect and delayed effects that complicate credit assignment, and (2) multi-dimensionality, where behaviors such as rapport‑building or knowledge‑seeking contribute indirectly to goal achievement. These characteristics make Markov decision process (MDP)-based RL with single-dimensional episode-level rewards inefficient and unstable. To address these challenges, we propose Sotopia-RL, a novel framework that refines coarse episode-level feedback into utterance-level, multi-dimensional rewards. Utterance-level credit assignment mitigates partial observability by attributing outcomes to individual utterances, while multi-dimensional rewards capture the full richness of social interactions and reduce reward hacking. Experiments in Sotopia, an open‑ended social learning environment, demonstrate that Sotopia-RL achieves state‑of‑the‑art social goal completion scores (7.17 on Sotopia-hard and 8.31 on Sotopia-full), significantly outperforming existing approaches. Ablation studies confirm the necessity of both utterance‑level credit assignment and multi‑dimensional reward design for RL training. Our implementation is publicly available at: https://github.com/sotopia-lab/sotopia-rl.
Figure 1: Sotopia-RL enhances social intelligence across diverse tasks. Sotopia-RL outperforms Sotopia-$\pi$ in goal completion score on accommodation, persuasion, collaboration, and negotiation tasks, as evaluated by Sotopia-eval. Sotopia-$\pi$ only utilizes episode-level goal completion score as reward signals to conduct self-reinforcement, while Sotopia-RL utilizes more fine-grained reward designs for RL training.
@@ 1 Introduction
Social intelligence (Gweon et al., 2023; Mathur et al., 2024; Zhu et al., 2025) has emerged as a critical capability for large language models (LLMs), enabling them to interpret and respond effectively to human feedback in multi-turn interactions (Han et al., 2024; Yi et al., 2024). It underpins a wide range of impactful applications, including customer service (Pandya & Holia, 2023; Bamberger et al., 2023), educational tutoring (Stamper et al., 2024; Nye et al., 2023), conflict resolution (Aggrawal & Magana, 2024), and team coordination (Li et al., 2023; Guo et al., 2024). In these interactions, communication proceeds through a sequence of utterances. We define an utterance as a discrete unit of communication that the agent deliberately produces as an action to pursue its social goals. Utterances are not merely linguistic tokens; they carry intentions, emotions, and shape the trajectory of the conversation. A socially intelligent agent must therefore generate high‑quality utterances that balance two aspects of effectiveness. At the single‑step level, an utterance should be goal‑related, directly advancing the agent’s social goal. Across multiple steps, it should also promote dialogue productivity, sustaining an engaging and constructive conversation that indirectly supports long‑term goals. This raises a central question for social learning: How can we define a high‑quality utterance and enable agents to generate them effectively?
@@
Uniqueness of social intelligence tasks. Unlike math and coding tasks commonly addressed by training LLMs with RL (Shao et al., 2024b; Wei et al., 2025; Cui et al., 2025), social tasks (like persuasion and collaboration) pose fundamentally different challenges. First, social interactions are inherently partially observable: agents can access the dialogue history but not latent factors such as intentions, emotions, or social norms that shape the outcome. As a result, social episodes are noisy, with the quality of individual utterances often only loosely correlated with the final success. In contrast, tasks like math or coding provide much clearer signals: a correct trajectory typically requires that most steps are high‑quality, with single‑step quality closely aligned to the final result. Second, social interactions are inherently multi‑dimensional: while some utterances contribute directly to goal achievement, others support it indirectly by maintaining rapport, fostering engagement, and sustaining conversational flow. Unlike math and coding tasks, where the outcome is mostly verifiable and binary (Su et al., 2025), social success must be analyzed across multiple dimensions.
Key challenges for training social agents. RL is a natural choice for social agent training because environments like Sotopia enable dynamic generation of multi‑turn social interactions, allowing agents to learn directly from feedback (Wang et al., 2024b; Ouyang et al., 2022). Unlike supervised learning, which depends on static datasets, RL captures the adaptive and evolving nature of dialogue. Effective RL training requires an online reward model (RM) to evaluate the quality of an utterance, and the unique features of social interactions make RM training difficult. Partial observability produces noisy, high‑variance episode‑level rewards, where low-quality utterances in the middle of the episode can lead to successful outcomes. We address this with utterance‑level credit assignment, which supplies fine‑grained supervision. At the same time, the multi‑dimensionality of social interactions risks RM overfitting to spurious features. To mitigate this, we incorporate additional reward dimensions—such as rapport‑building and knowledge‑seeking—that regularize learning and align the RM with broader social expectations.
Key insights. Our key insights are that, despite the complexity of social interactions, reward assignment for social interaction is both feasible and reliable when powered by LLMs: (1) Episode‑level outcomes can be decomposed into utterance‑level contributions, and LLMs can reliably infer which utterances contributed to or hindered success. Empirically, diverse LLM families (e.g., Gemini‑2.5‑flash (Google DeepMind, 2024), GPT‑4o (Hurst et al., 2024), Qwen2.5‑7b (Qwen et al., 2025)) show strong agreement in attribution scores, with Spearman correlations exceeding 0.7. (2) Evaluating utterance quality across distinct sub‑dimensions—such as rapport, engagement, and knowledge‑sharing—simplifies the credit assignment task, turning a difficult holistic judgment into a set of rubric‑based evaluations and making the reward design with less variance. Together, these insights show why utterance‑level and multi‑dimensional reward design is practical without human annotation, and why LLMs align well with human preferences.
@@
Main discoveries. Based on these key insights, we utilize Sotopia (Zhou et al., 2023a), an open-ended social learning environment, for the RL-based training of social agents. Sotopia provides diverse social scenarios for simulation and multi-dimensional evaluation metrics named as Sotopia-eval. Experiments conducted in the Sotopia environment reveal two key findings: (1) Social agents trained with Sotopia-RL consistently outperform all baselines on social goal completion metrics evaluated by Sotopia-eval, achieving a goal completion score of 7.17 on the Sotopia-hard benchmark and 8.31 on the full Sotopia dataset. (2) Both utterance-level reward attribution and multi-dimensional reward combination for reward design are critical for stable and effective RL training in complex social scenarios. These results highlight the importance of social reward design and validate the core design principles behind Sotopia-eval —particularly the importance of evaluating social interaction quality across diverse dimensions.
@@ 2 Related Work
Social skill learning. To augment the social intelligence of agents, prior work utilizes RL to achieve this. Sotopia-$\pi$ (Wang et al., 2024b) uses self-reinforcement learning, Ndousse et al. (2021) employ conversation-level RL rewards, and Stable Alignment (Pang et al., 2024) trains via rule-based peer feedback without rewards. SDPO (Kong et al., 2025) uses preference-based tuning but overlooks utterance-level impact. Our key contribution is the design of utterance-level rewards specifically for social tasks. Instead of conducting explicit strategy injection for training (Zhang et al., 2025; Wang et al., 2025), we implicitly model social skills within the design of our rewards.
Process reward modeling. For reinforcement learning with verifiable rewards (RLVR) like math and programming, Process Reward Modelings (PRMs) have been effectively utilized. PRIME (Cui et al., 2025) assigns token-level rewards using only outcome labels, enhancing reasoning capabilities without the need for explicit process annotations. Similarly, other works (Choudhury, 2025; Wang et al., 2024a) employ Monte Carlo (MC) rollouts to compute reward targets, enabling scalable training of RL. MC methods estimate the expected value of uncertain quantities through repeated random sampling, making them particularly effective in modeling stochastic processes and complex static decision-making tasks (Barto, 2021). Designing an utterance-level reward for social tasks can also be considered as one special type of process reward. Due to the ambiguity and multi-dimensionality of social tasks, such process reward requires multi-dimensional evaluation.
Multi-objective RL. Multi-objective RL provides a foundational approach for aligning LLMs with multiple preferences by framing alignment as a multi-objective optimization problem. The most common strategy employs linear scalarization, which combines multiple reward functions into a single objective (Jang et al., 2023; Yang et al., 2024; Li et al., 2020; Zhou et al., 2023b), enabling the reuse of standard RL techniques while varying reward weights. Recent work extends this to non-linear combinations by incorporating utility functions (Cheng et al., 2025) or using LLMs to search for reward functions (Xie et al., 2024). Other studies improve reward modeling by decomposing rewards into more informative components (Mao et al., 2025; Shenfeld et al., 2025; Lee et al., 2024). Building on these directions, our work focuses on multi-dimensional reward learning for social tasks and relies on linear scalarization, where auxiliary rewards—such as relationship maintenance and knowledge seeking—are explicitly modeled to support goal completion.
3 Sotopia Environment
In this section, we formally define the Sotopia environment, covering social interactions along with their observation and action spaces, as well as social evaluations. This definition highlights the partial observability and multi‑dimensional nature of social interactions.
@@ 3.1 Social Interaction
Given the set $I=\{A,B\}$ of participants in a dialogue, social interaction can be modeled as a partially observable Markov decision process (POMDP), represented by the tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\right\rangle$. $\mathcal{S}$ represents the set of possible social states, $\mathcal{A}$ represents the action space, $\mathcal{O}$ represents the observation space. $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. $Z:\mathcal{S}\rightarrow\mathcal{O}$ is the observation function. $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that reflects the overall quality of the social interaction with respect to each agent’s private goal, such as successful persuasion or mutual understanding.
Observation space. In Sotopia, the observations refer to the history of dialogue. Each agent $i\in I$ operates under partial observability, receiving at each time step a private observation $o_{t}^{i}\in\mathcal{O}_{i}$, where $\mathcal{O}_{i}\subseteq\mathcal{O}$ denotes the private observation space of agent $i$. $\mathcal{O}_{i}$ consists of all dialogue histories and contextual cues that agent $i$ can perceive, but excludes latent variables such as the partner’s private goals, beliefs, or emotions. The observation $o_{t}^{i}$ is generated by the probabilistic observation function $Z$, modeling the partial and asymmetric nature of social perception. A social episode with $T$ turns observed by agent $i$ is defined as
$\tau_{i}=\big{(}o_{0}^{i},a_{0}^{i},;o_{1}^{i},a_{1}^{i},;\ldots,;o_{T}^{i}\big{)},$ (1)
where $o_{t}^{i}\in\mathcal{O}_{i}$ is the dialogue history observed at time $t$, and $a_{t}^{i}\in\mathcal{A}_{i}$ is the utterance generated by agent $i$ at time $t$. This episode captures the full sequence of observations and actions.
Action space. An action in Sotopia is defined as an utterance, including both verbal (e.g., speaking) and non‑verbal (e.g., smiling, hugging) communication. For each agent $i$, the action space $\mathcal{A}_{i}$ contains all such communicative behaviors. At turn $t$, the agent samples $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid o_{t}^{i},g_{i})$ from its policy conditioned on observation $o_{t}^{i}$ and private goal $g_{i}$. The joint actions then contribute to and become part of the next observation $o_{t+1}^{i}$.
MDP approximation. To optimize the social agent with MDP-based RL methods like GRPO (Shao et al., 2024b), we adopt an MDP approximation of the Sotopia environment. At step $t$, the state $s_{t}^{i}$ is considered as the dialogue history, together with its private social goal, the social policy $\pi_{\theta}$ outputs a distribution over utterances $a_{t}^{i}$ (i.e., $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid s_{t}^{i})$), and offline rewards $r_{t}^{i}$ from Stage 1 (Figure 4) are used to train a online utterance-level reward model $R_{\psi}(s_{t}^{i},a_{t}^{i})$.
@@ 3.2 Social Evaluation
To support RL-based training, the Sotopia environment provides quantitative feedback on how well each agent achieves its goal using Sotopia-eval, a multi-dimensional evaluation suite with an LLM-based evaluator $f_{\theta}$.
Multi-dimensional evaluation.
For each agent $i$, the LLM-based evaluator produces a score conditioned on the social episode $\tau_{i}$ and the agent’s goal $g_{i}$: $G_{i}=f_{\theta}(\tau_{i},g_{i})\in\mathbb{R}$. This score represents the extent to which the agent $i$ achieves its social goal $g_{i}$ through social interactions $\tau_{i}$. In Sotopia-eval, this single score is extended to a seven-dimensional vector $\mathbf{G}_{i}\in\mathbb{R}^{7}$, expanding from goal completion (Goal) to believability (Bel), knowledge seeking (Kno), relationship maintaining (Rel), secret keeping (Sec), social rule-following (Soc), and financial/material benefits (Fin). Among these, Goal serves as the primary objective, while the others provide auxiliary signals of social interaction quality. Detailed definition for each dimension is available in Appendix §H.

### v3

@@ Abstract
Social intelligence has become a critical capability for large language models (LLMs), enabling them to engage effectively in real-world social tasks such as collaboration and negotiation. Reinforcement learning (RL) is a natural fit for training socially intelligent agents because it allows models to learn sophisticated strategies directly through social interactions without requiring human annotations. However, there are two unique parts about social intelligence tasks: (1) the quality of individual utterances in social interactions is not strictly related to final success; (2) social interactions require multi-dimensional rubrics for success. Therefore, we argue that it is necessary to design rewards for building utterance-level multi-dimensional reward models to facilitate RL training for social intelligence tasks. To address these challenges, we propose Sotopia-RL, a novel framework that refines coarse episode-level feedback into utterance-level, multi-dimensional rewards. Utterance-level credit assignment attributes outcomes to individual utterances, while multi-dimensional rewards capture the full richness of social interactions and reduce reward hacking. Experiments in Sotopia, an open‑ended social learning environment, demonstrate that Sotopia-RL achieves state‑of‑the‑art social goal completion scores (7.17 on Sotopia-hard and 8.31 on Sotopia-full), significantly outperforming existing approaches. Ablation studies confirm the necessity of both utterance‑level credit assignment and multi‑dimensional reward design for RL training.
@@ 1 Introduction
Social intelligence (Gweon et al., 2023; Mathur et al., 2024; Zhu et al., 2025), a capability with applications in customer service (Pandya & Holia, 2023; Bamberger et al., 2023), educational tutoring (Stamper et al., 2024; Nye et al., 2023), conflict resolution (Aggrawal & Magana, 2024), and team coordination (Li et al., 2023; Guo et al., 2024), has emerged as a crucial capability for large language models (LLMs). Social intelligence is naturally formed via social interactions, and Reinforcement learning (RL) is a natural fit for training socially intelligent agents because it allows models to learn sophisticated strategies directly through social interactions. Therefore, utilizing reinforcement learning (RL) for social intelligence learning is natural. Although prior work has shown that RLs can be used to optimize LLM abilities for math (Shao et al., 2024b; Yang et al., 2024a), coding (Hui et al., 2024; Wei et al., 2025), and reasoning (Guo et al., 2025), tuning mainstream LLMs (Yang et al., 2025; Touvron et al., 2023; Hurst et al., 2024) for social intelligence with RL remain underexplored (Zhou et al., 2025; Mathur et al., 2024). We argue that a central obstacle is the lack of a unified and effective reward design that can be tailored to social behavior and be applied to diverse types of social scenarios like collaboration, negotiation and accommodation. In this paper, we propose a practical reward-design recipe for social intelligence and apply reinforcement learning to train social agents that generate high-quality social utterances.
@@
Uniqueness of social intelligence tasks. Unlike math and coding tasks, which often provide clear and verifiable rewards (Shao et al., 2024b; Wei et al., 2025; Cui et al., 2025), social tasks (e.g., persuasion and collaboration) present fundamentally different challenges for reward design. First, the quality of individual utterances in social interactions is hard to define and often only loosely correlates with the final success: in a negotiation, for example, a social agent might deploy a misleading claim that nevertheless helps secure a better outcome. By contrast, domains such as math or coding give much clearer signals, where correct intermediate steps are usually required for a correct outcome. Second, social interactions are inherently multi-dimensional: some utterances in the social interactions directly advance the task goal, while others play an indirect but crucial role by building rapport, maintaining engagement, or preserving conversational flow. Unlike largely verifiable, often binary outcomes in math or coding (Su et al., 2025), social interactions must be considered and evaluated across multiple interacting dimensions.
Necessity of utterance-level multi-dimensional reward models for social intelligence tasks. Training social agents with only episode outcomes is common in principle but is highly sample-inefficient and often hard to capture fine-grained conversational behaviors. It is mainly because utterance-level contributions weakly correlate with episode outcomes. Instead, we argue for utterance-level, multi-dimensional reward models (RM): first, episode outcomes can be decomposed into utterance-level contributions, and LLMs are capable of reliably inferring which utterances helped or hindered success—empirically, diverse LLM families show strong agreement in attribution (with Spearman correlations >> 0.7). Second, evaluating utterance quality along separate sub-dimensions (e.g., goal achieving, relationship building, knowledge-sharing) builds a rubric-based assessment for difficult holistic judgement, which simplifies credit assignment and reduces variance in reward estimates. Together, these observations explain why multi-dimensional, utterance-level reward design is both practical—often without costly human annotation—and better aligned with human preferences than monolithic episode-level ones (Ram’e et al., 2024; Moskovitz et al., 2023).
Our method. In this paper, we propose Sotopia-RL, a novel and practical RL training receipt for social agent. It includes two main stages: (1) offline social reward collection; (2) online social agent training. For the first stage, we collect a set of existing social interaction episodes with outcomes and utilize LLMs and multi-dimensional rubrics to conduct utterance-level, multi-dimensional social reward assignment. For the second stage, we utilize such offline-collected social reward to train an utterance-level, multi-dimensional RM and conduct online RL training.
Figure 1: Uniqueness of social intelligence tasks. (Left) Accommodation, persuasion, collaboration, and negotiation represent four core types of social intelligence tasks. Our method achieves consistent improvements across all tasks compared with Sotopia-$\pi$, the previous state-of-the-art on Sotopia. (Right) Two unique features of social interactions: (1) utterances are not always directly tied to outcomes—e.g., people may lie or mislead in negotiations; (2) interactions are inherently multi-dimensional—e.g., relationship building can play a crucial role in achieving task success.
@@
Main discoveries. To prove the effectiveness of our proposed RL methods, we evaluate based on Sotopia (Zhou et al., 2023a), an open-ended social learning environment. Sotopia provides diverse social tasks and multi-dimensional social evaluation. Experiments conducted in the Sotopia environment reveal two key findings: (1) Social agents trained with Sotopia-RL consistently outperform all baselines on social goal completion metrics provided by Sotopia, achieving a goal completion score of 7.17 on the Sotopia-hard benchmark and 8.31 on the full Sotopia dataset. (2) Our utterance-level and multi-dimensional reward design is critical for stable and effective RL training in complex social scenarios. These results highlight the importance of social reward design and validate the core design principles behind Sotopia —particularly the importance of evaluating social interaction quality across diverse dimensions.
@@ 2 Related Work
Utterance in social interactions. A social utterance is a deliberate communicative act produced to pursue social goals (Bahing et al., 2018). Social interactions are sequences of such utterances, with multi-turn exchanges forming an episode. Beyond linguistic tokens, utterances convey intentions and emotions that shape the conversation (Ghosal et al., 2020; Wang et al., 2022).
Social skill learning. To enhance agents’ social intelligence, prior approaches have leveraged RL in different ways. Sotopia-$\pi$ (Wang et al., 2024b) adopts self-reinforcement learning, Ndousse et al. (2021) use conversation-level rewards, and Stable Alignment (Pang et al., 2024) relies on rule-based peer feedback without explicit rewards. SDPO (Kong et al., 2025) incorporates preference-based tuning but ignores utterance-level effects. Our contribution is to introduce utterance-level reward modeling tailored for social tasks. Rather than injecting explicit strategies during training (Zhang et al., 2025; Wang et al., 2025), we capture social skills implicitly through the reward design.
Process reward modeling. For RL with verifiable rewards (RLVR) such as math and programming, Process Reward Models (PRMs) have proven effective. PRIME (Cui et al., 2025) assigns token-level rewards from outcome labels, boosting reasoning without explicit process annotations. Other works (Choudhury, 2025; Wang et al., 2024a) use Monte Carlo (MC) rollouts to compute reward targets, where repeated sampling estimates expected values in stochastic or complex decision-making tasks (Barto, 2021). Designing utterance-level rewards for social tasks can be viewed as a form of process reward, but unlike math or programming, the ambiguity and multi-dimensionality of social interactions demand multi-dimensional evaluation.
Multi-objective RL. Multi-objective RL aligns LLMs with multiple preferences by optimizing over several reward functions. Most approaches use linear scalarization to combine rewards into a single objective (Jang et al., 2023; Yang et al., 2024b; Li et al., 2020; Zhou et al., 2023b), while recent work explores non-linear utilities (Cheng et al., 2025; Xie et al., 2024) and reward decomposition (Mao et al., 2025; Shenfeld et al., 2025; Lee et al., 2024). Building on these directions, we focus on multi-dimensional reward learning for social tasks, using linear scalarization where auxiliary rewards (e.g., relationship maintenance, knowledge seeking) explicitly support goal completion.
3 Social Learning Environment
To enable RL training, we first need to define a suitable environment. Such an RL environment must serve two purposes: (i) it should generate data by simulating social interactions, and (ii) it should provide feedback that can be used as training signals. Following Sotopia (Zhou et al., 2023a), we implement both aspects: interaction trajectories are generated through multi-turn dialogues between agents, and episode-level feedback is provided through LLM-based evaluation.
@@ 3.1 Social Interaction
Social interactions are dialogues between a pair of agents. From the perspective of a single agent, another side of the social agent is considered as the partner model. The interaction process can be described as a partially observable Markov decision process (POMDP), represented by the tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\right\rangle$. $\mathcal{S}$ represents the set of possible social states, $\mathcal{A}$ represents the action space, $\mathcal{O}$ represents the observation space. $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function that captures how the social state evolves given the agent’s utterance and the partner’s response. $Z:\mathcal{S}\rightarrow\mathcal{O}$ is the observation function. $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that reflects the overall quality of the social interaction with respect to the agent’s private goal, such as successful persuasion or mutual understanding.
Observation space. In a social learning environment like Sotopia, the observations refer to the history of dialogue. The social agent operates under partial observability, receiving at each time step $t$ a private observation $o_{t}\in\mathcal{O}$. $\mathcal{O}$ consists of all dialogue histories and contextual cues that the agent can perceive, but excludes latent variables such as the partner’s private goals, beliefs, or emotions. The observation $o_{t}$ is generated by the probabilistic observation function $Z$, modeling the partial and asymmetric nature of social perception. A social episode with $T$ turns is defined as
$\tau=\big(o_{0},a_{0},\,o_{1},a_{1},\,\ldots,\,o_{T}\big),$ (1)
where $o_{t}\in\mathcal{O}$ is the dialogue history observed at time $t$, and $a_{t}\in\mathcal{A}$ is the utterance generated by the agent at time $t$. This episode captures the full sequence of observations and actions.
Action space. An action in Sotopia is defined as an utterance, including both verbal (e.g., speaking) and non-verbal (e.g., smiling, hugging) communication. The action space $\mathcal{A}$ contains all such communicative behaviors available to the agent. At turn $t$, the agent samples $a_{t}\sim\pi_{\theta}(\cdot\mid o_{t},g)$ from its policy conditioned on the current observation $o_{t}$ and its private goal $g$. The chosen action, together with the partner’s response, contributes to and becomes part of the next observation $o_{t+1}$.
MDP approximation. To optimize the social agent with MDP-based RL methods like GRPO (Shao et al., 2024b), we adopt an MDP approximation. At step $t$, the state $s_{t}$ is considered as the dialogue history together with the agent’s private social goal. The social policy $\pi_{\theta}$ outputs a distribution over utterances $a_{t}$ (i.e., $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})$), and offline rewards $r_{t}$ are used to train an online utterance-level reward model $R_{\psi}(s_{t},a_{t})$.
@@ 3.2 Social Evaluation
As a social learning environment for RL training, it should be able to provide quantitative feedback on the outcome of the social interaction. Therefore, we adopt the multi-dimensional social evaluation in Sotopia as environment feedback. In particular, we define the goal completion score $G$ as the primary training signal, obtained from an LLM-based evaluator $f_{\theta}$ conditioned on the social episode $\tau$ and the agent’s private goal $g$:
$G=f_{\theta}(\tau,g)\in\mathbb{R}.$ (2)
Beyond goal completion, it also produces six auxiliary evaluation dimensions—believability, knowledge seeking, relationship maintaining, secret keeping, social rule-following, and financial/material benefits—following the Sotopia rubric. These additional dimensions are carefully defined for analysis and provide richer insights into the quality of social interactions. Detailed definitions for each dimension are given in Appendix §I.